Fermi National Laboratory


Accelerator Update

Monday December 29
The day shift began with the Tevatron (TeV) and the Antiproton Source (Pbar) in access, and with TeV sector B1 cooldown progressing.

At 9:47 AM, Linac technicians accessed their tunnel. By 1 PM, the top slot of tanks 2 through 5 had been cleaned out and the upstream quadrupole of tank five, which had developed a ground fault, had been bypassed.

Operations established beam to MiniBooNE at 3 PM.

Recycler experts began a study period at 4:39 PM.

At 8:22 PM, Operations sent beam to Pbar. Stacking began at 8:40 PM.

The TeV suffered a sector B1 quench at 8:49 PM.

Tuesday December 30
The midnight shift began with the TeV in start up, with Pbar stacking, with MiniBooNE taking beam, and with Main Injector (MI) and Switchyard (SY) experts conducting 120 GeV slow extraction studies.

At 12:19 AM, Recycler experts completed their study.

The TeV suffered another B1 quench at 1:33 AM. At 2:40 AM, experts accessed the TeV to check lead flows. They were out of access by 3:04 AM. Everything looked okay. TeV experts began a study period.

At 10:26 AM, due to an error with the shot lattice, most of the Pbar stack was lost.

SY studies ended at 6:19 PM.

At 9:12 PM, TeV experts established a 36x0 study-store. They then established a 14x0 store to conduct an abort study.

Operations put the TeV into shot setup at 10:10 PM.

Wednesday December 31
The midnight shift began with the TeV in shot setup, with some problems in Pbar, and with MiniBooNE taking beam.

At 2:19 AM, Operations established store 3118 with an initial luminosity of 15.24E30. Pbar resumed stacking.

Recycler experts began a study period at 4 AM.

Around 8 AM, Operations received an indication of a Booster LCW leak. After checking everywhere else, operators made Booster ready for an access to check the tunnel at 9:45 AM. At 10:18 AM, operators found a magnet hose leak. The leak was repaired by noon.

Operations terminated store 3118 at 12:04 PM.

Stacking resumed at 12:18 PM.

D-zero (D0) experimenters began a thirty-minute access to their collision hall at 12:23 PM.

Beam to MiniBooNE resumed at 12:26 PM.

Operations put the TeV through a dry squeeze at 1:13 PM and into shot setup at 1:44 PM.

At 4:19 PM, Operations established store 3120 with an initial luminosity of 27.83E30. Pbar resumed stacking.

Thursday January 1 - HAPPY NEW YEAR!
The midnight shift began with Operations monitoring store 3120, with Pbar stacking, and with MininBooNE taking beam.

Operations terminated store 3120 at 10 AM and then put the TeV into shot setup a little while later.

At 12:35 PM, Operations established store 3122 with an initial luminosity of 44.73E30. Pbar resumed stacking.

The stacking rate had been very poor. Operations called for expert help at 7 PM. That help arrive at 8:22 PM.

Friday January 2
The midnight shift began with Operations monitoring store 3122, with Pbar stacking, and with MiniBooNE taking beam.

MI and SY experts began slow extraction studies, and Recycler experts began reverse proton studies at 8 AM. These studies ran for three hours.

At 8:33 AM, The TeV suffered a sector A3 quench.

At 9:27 AM, all the 750 KeV Line power supplies tripped off. Operators investigated and discovered a Linac water skid off. An expert found and bypassed a bad flow meter and had the system back up by 10:22 AM.

D0 experimenters access their collision hall at 9:51 AM. They were out of access thirty minutes later.

Operations put the TeV into shot setup at 11:30 AM.

Operators made Booster ready for access at 3:44 PM so techs could repair a RF cavity (BRF17). A Linac expert began replacing the bad flow meter in the 750 KeV Line.

At 4:16 PM, Operations established store 3123 with an initial luminosity of 44.3E30.

A Linac RF station (KRF1) tripped off. Operators reset the station 4:39 PM.

Pbar resumed stacking at 8:36 PM and MiniBooNE resumed tacking beam at 8:43 PM.

Saturday January 3
The midnight shift began with Operations monitoring store 3123, with Pbar stacking, and with MiniBooNE taking beam.

Operations contacted a Linac expert at 4 AM. He reported that the transformer that serves KRF1 was tripping off due to high humidity and needs its oil processed. A tech will be called in during the day to do this.

At 10:12 AM, operators discovered a LCW leak on KRF1 and called techs to make repairs. This work was completed at 12:15 PM.

Among other Pbar problems, Operations noticed at 1:14 PM that some Pbar diagnostics were not working properly. Operators called in experts. By 2:41 PM, the Accumulator was getting small amounts of beam.

At 6 PM, experts began the oil processing for KRF1.

At 7:19 PM, Pbar experts discovered a problem with a Septa (A:ISEP2V) that required an access. This will occur after the next shot.

Operations terminated store 3123 at 8:15 PM.

At 9:25 PM, Operation put the TeV into shot setup.

Operations established store 3125 at 11:14 PM with an initial luminosity of 43.1E30. Pbar resumed stacking.

At 11:32 PM, operators prepared Pbar for access.

Sunday January 4
The midnight shift began with Operations monitoring store 3125, with experts accessing Pbar, and with MiniBooNE taking beam.

At 12:50 AM, the Pbar access had ended and Operations began stacking. An expert found and tightened a loose connection.

MiniBooNE suffered from some beam drifting at 1:52 PM.

At 3:15 PM, Linac suffered a 400 MeV area trip. Operators discovered that the I- Source tripped off. They reset the source, but it tripped off several more times. Operators investigated for problems, but found none. By 3:48 PM, they had it working normally while in local mode. It was put back in remote at 7 PM.

At 4:49 PM, all the MI RF stations tripped off due to an errant command. Everything reset and came back on okay.

MiniBooNE suffered three trips between 7 PM and 7:42 PM.

Monday January 5
The midnight shift began with Operations monitoring store 3125, with Pbar stacking, and with MiniBooNE taking beam.

At 12:46 AM, Pbar problems caused the antiprotons to begin unstacking. By 2:20 AM the beam was recovering slowly. An expert thought the problem was due to the large stack. About 9 mA were lost from the stack.

Operations terminated store 3125 at 2:25 AM.

Operations put the TeV into shot setup at 3:02 AM. At 4:39 AM, Operations established store 3127 with an initial luminosity of 45.62E30. Pbar resumed stacking.

Today's Plans
The plan for today is to allow Pbar experts to work on stacktail problems and to continue to stack and store.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 1/5/2004   email Fermilab